OpenStudy (anonymous):

Exactly 100 mL of .10 M HNO2 are titrated with a .10 M NaOH solution. What is the pH at the equivalence point? Ka (HNO2) = 5.6*10^-4

OpenStudy (anonymous):

Could someone provide the reactions as well?

OpenStudy (australopithecus):

pH = -log([H+])

OpenStudy (australopithecus):

HNO2 + NaOH -> H2O + NaNO2

OpenStudy (australopithecus):

notice it is a 1 to 1 reaction

OpenStudy (anonymous):

Do you need to react NaNO2 with H2O ?

OpenStudy (australopithecus):

Ka = [A-][H+]/[HA]

OpenStudy (australopithecus):

https://en.wikipedia.org/wiki/Acid_dissociation_constant

OpenStudy (australopithecus):

ok to the actual problem

OpenStudy (australopithecus):

https://en.wikipedia.org/wiki/Equivalence_point

OpenStudy (australopithecus):

http://www.chemguide.co.uk/physical/acidbaseeqia/sasb2.gif

OpenStudy (australopithecus):

that link will explain to you what the equivalence point is, it is the point in which 50%% of the acid is protonated and 50% of the acid is deprotonated

OpenStudy (australopithecus):

You can solve this problem by simply finding the pKa

OpenStudy (australopithecus):

pKa is the point at which a chemical is 50% protonated and 50% deprotonated

OpenStudy (anonymous):

Well. The reaction you provided produced a basic salt. So that product should react with water next, in order to produce OH-. NO2- + H2O -- HNO2 + OH-. So I can use the Kb to find [OH-]. And from there I can find pOH and subtract it from 14 to get pH.

OpenStudy (anonymous):

I got it now. Thank you!

OpenStudy (australopithecus):

yeah Im rusty with this stuff I'm not going to lie and it is 4am here, sorry for not providing you with a very sound answer, im much too tired to be coherent

OpenStudy (anonymous):

Haha no worries. I'm just cramming for an exam tomorrow. Have a good night!

OpenStudy (australopithecus):

\[pKa = -\log(5.6*10^{-4})\] Pka = 7.48757 pH = pKa + log[A/HA] A = 0.1M Ha = 0.1M Log[0.1/0.1] = 0 therefore, pH = pKa like I state earlier

OpenStudy (australopithecus):

stated*

OpenStudy (australopithecus):

This is because the reaction is 1 to 1 as you should see above therefore the equivalence point is when the concentration of the conjugate base is equal to the concentration of the acid

OpenStudy (australopithecus):

I gave you the answer but I failed to prove to you why pKa = pH at which an acid is at its equivalence point

OpenStudy (australopithecus):

Since, pH = pKa Pka = -log(Ka) therefore pH = -log(Ka) at the equivalence point

OpenStudy (australopithecus):

Therefore the pH at the equivalence point is 7.48757

OpenStudy (australopithecus):

sorry for ignoring sig figs